# Continual Learning on CLIP via Incremental Prompt Tuning with Intrinsic Textual Anchors

### v2

@@ 3.2 Overview
We present the framework of our methods in Fig. 3. Previous CLIP-based CL methods (Wang et al., 2022c; 2023; Zhou et al., 2023c) usually treat the text encoder as a pre-trained classifier, overlooking the rich representation capability of the language modality. In Sec. 3.3 we introduce TPPT-V, which guides the learning of incremental visual prompts with static textual prototypes. Since learning with static textual prototypes may be limited by the quality of prototypes, in Sec. 3.4 we further propose TPPT-VT to learn textual prompts for higher quality and representative textual prototypes. Our approach encourages textual diversity to better regulate the learning process.
@@ 3.3 Incremental Visual Prompt Tuning with Textual Prototypes
Incremental visual prompt tuning. Prompt-based methods in CL have been proven to be highly effective in mitigating catastrophic forgetting, primarily by utilizing a fixed or incremental set of prompts (Wang et al., 2022e; d; Smith et al., 2023). Building on this, we employ a straightforward method to prepend an incremental set of visual prompts on multiple Transformer layers (Jia et al., 2022; Rasheed et al., 2023; Khattak et al., 2023b). These prompts are denoted as ${\mathbf{P}}_{v,l}=\{{\mathbf{P}}_{v,l}^{1},{\mathbf{P}}_{v,l}^{2},\cdots,{\mathbf{P}}_{v,l}^{M}\}$ for each Transformer layer $l$ ranging from the first layer up to the $d_{v}$-th layer of the vision encoder, where $M$ is the total number of prompts, and $d_{v}$ is a hyperparameter indicating the depth of Transformer layers at which prompts are learned.
@@
${\mathbf{P}}_{v,l}=\sum_{m=1}^{M}\alpha_{m}{\mathbf{P}}_{v,l}^{m},$ (2)
where the weights $\alpha=\text{LN}(q({\mathbf{x}}))$ are obtained from a simple linear layer applied to the queried input $q({\mathbf{x}})$, in contrast to the complicated attended query-key matching with additional learnable attention vectors used in (Smith et al., 2023). Here, the query function $q(\cdot)=f_{\theta}(\cdot)\in\mathbb{R}^{D}$ refers to the pre-trained image encoder (Wang et al., 2022e; d; Smith et al., 2023). By leveraging the powerful language modality, our formulation is simple yet effective, eliminating the need for complex incremental learnable attention vectors and orthogonal regularization in (Smith et al., 2023). The prompt tuning process is guided by consistent and stable knowledge derived from previously learned information through textual prototypes during incremental stages.
Anchoring the past with textual prototypes. Existing CLIP-based CIL methods (Wang et al., 2023; Thengane et al., 2022) often utilize the text embeddings merely as a pre-trained classifier, thereby not fully exploiting the rich representational capacity of the language modality. To this end, we propose TPPT-V to guide the learning of incremental visual prompts through the use of static textual prototypes as anchor points in the embedding space. TPPT-V mitigates catastrophic forgetting by consistently ensuring that visual representations remain closely aligned and compact around corresponding static textual prototypes.
Specifically, our textual prototypes are derived from text embeddings, obtained through hand-crafted prompt templates with class names from OpenAI [^1] https://github.com/openai/CLIP whenever available. For datasets where prompt templates are not provided, we apply the basic template of “a photo of a [cls]”. To enhance the alignment (Wang & Isola, 2020) between visual features and their corresponding textual prototypes, we propose the Textual Prototypical Contrastive Loss (Li et al., 2020; Cui et al., 2023; Li et al., 2023). This approach leverages the robust pre-trained representational capacity of these textual prototypes while preventing visual features from being assigned to the classes previously learned. It also effectively prevents the drift of previously learned semantics throughout the incremental learning phase (Fig. LABEL:fig:drift). We denote the Textual Prototypical Contrastive Loss as:
@@
where $\mathbbm{1}(\cdot)$ is the indicator function that selects samples corresponding to the given textual prototype. $\tilde{{\mathbf{z}}}=\tilde{f_{\theta}}({\mathbf{x}})$ is the visual embedding obtained through the prompted image encoder, and ${\mathbf{w}}_{c}$ is the text embedding for class $c$, acting as a static textual prototype. $C$ is the number of classes already seen.
@@
where $\mathcal{L}_{\text{CE}}$ represents the cross-entropy loss between predicted probabilities and ground truth. All loss terms are computed with the prompted image encoder $\tilde{f_{\theta}}$ and the frozen pre-trained text encoder $g_{\psi}$. This loss shares a similar form with the pre-training loss of the CLIP model (Radford et al., 2021), but differs in purpose. While CLIP focuses on the one-to-one mapping between an image and its own caption to form a broad representation space, our method aims to cluster image embeddings to be compact around their corresponding class text embeddings, from the perspective of CL to mitigate forgetting.
@@ 3.4 Incremental multi-modal prompt tuning with regularized textual prototypes
Incremental textual prompt tuning. Although static textural prototypes from hand-crafted text prompts of classes effectively serve as stable anchors in TPPT-V, the effectiveness can be limited by the gap between pre-trained CLIP and specific downstream tasks (e.g., fine-grained classification), as well as the inherent discrepancy between the textual and visual embeddings in CLIP (Liang et al., 2022). To address these limitations, we introduce TPPT-VT to further incorporate multi-model prompt tuning using learnable soft text prompts. Similar to the design of visual prompts in TPPT-V, we use an incremental set of textual prompts for each Transformer layer $l$, ranging from the first layer up to the $d_{t}$-th layer of the text encoder, where $d_{t}$ is set as a hyperparameter indicating the depth of Transformer layers at which prompts are learned. We denote these prompts as ${\mathbf{P}}_{t,l}=\{{\mathbf{P}}_{t,l}^{1},{\mathbf{P}}_{t,l}^{2},\cdots,{\mathbf{P}}_{t,l}^{C}\}$, where each class is assigned to a unique textual prompt, and $C$ represents the total number of classes encountered thus far. By using unique prompts for each class, TPPT-VT further enhances the fixed textual prototypes in TPPT-V, making them more representative.
Regularizing textual prompt tuning. While guiding the model with textual prototypes derived from hand-crafted prompts has demonstrated superior performance, the joint training of textual and visual prompts introduces new challenges. One critical issue is the model collapse, i.e, both branches of the joint-embedding model produce nearly identical and constant output vectors (Jing et al., 2021; Bardes et al., 2021) (as illustrated in Fig. LABEL:fig:div). This collapse compromises the model’s ability to distinguish between diverse inputs and effectively learn meaningful representations. In addition, naïve joint training can adversely affect the uniformity of the pre-trained model (Wang & Isola, 2020). Uniformity is a desirable property for joint-embedding models, ensuring that feature vectors are uniformly distributed roughly on the unit hypersphere. This uniform distribution is essential for preserving the maximum amount of information from the data.
To address these issues, we introduce the textual prototype diversity loss, which regularizes the model by explicitly enforcing textual prototypes to be uniformly distributed on the unit hypersphere. This is achieved by promoting high pairwise distances among textual prototypes (Wang & Isola, 2020; Cho et al., 2023), formally defined as follows:
@@ 4.1 Experiment Settings
Datasets. In line with common CIL benchmarks (Rebuffi et al., 2017; Wang et al., 2022e; d; Smith et al., 2023), we evaluate the performance using widely used datasets CIFAR100 (Krizhevsky et al., 2009), ImageNet-R (Deng et al., 2009), TinyImageNet (Le & Yang, 2015), and CUB200 (Wah et al., 2011). To further validate the efficacy of our approach, we extend our evaluation to include two additional datasets, FGVCAircraft (Maji et al., 2013) and Stanford Cars (Krause et al., 2013), following the popular VLM methods (Zhou et al., 2022c; Yao et al., 2023; Khattak et al., 2023a; b). Specifically, we split the training datasets into 10 tasks, resulting in 10 classes per task for CIFAR100, Aircraft, and Stanford Cars, 20 classes per task for ImageNet-R, TinyImageNet, and CUB200.
Training details. For a fair comparison, we use the same backbone model CLIP with ViT-B/16, using pre-trained weights from (Ilharco et al., 2021; Radford et al., 2021). We train all methods with a batch size of 64 for 5 epochs, using SGD with momentum for optimization. The learning rate starts from 0.1 and decays with cosine annealing. The visual prompt length $L_{v}$ and the textual prompt length $L_{t}$ are set to 4, and the depth for both visual prompts $d_{v}$ and textual prompts $d_{t}$ are set to 12. The number of prompts per task is set to 10. The loss weight $\alpha$ is set to 1 by default. Open source code will be made public upon acceptance.
Evaluation metrics. To demonstrate the effectiveness of our approach, we report two commonly used metrics (Rebuffi et al., 2017): (1) the final accuracy ${\mathcal{A}}_{T}$ after learning all the $T$ incremental stages, (2) the average accuracy $\bar{{\mathcal{A}}}=\frac{1}{T}\sum_{t=1}^{T}{\mathcal{A}}_{t}$ of the individual accuracies across the incremental stages.
4.2 Experimental Results
Comparisons with prompt-based methods. In Table 1, we evaluate the final and average accuracies across various datasets, comparing current state-of-the-art methods. Additionally, we illustrate the performance across incremental stages in Fig. LABEL:fig:exp_task, and provide evaluations of multiple independent runs with varied orders of the incoming classes in Appendix C. Our proposed methods, TPPT-V and TPPT-VT, exhibit a remarkable improvement over previous approaches. In contrast to previous prompt tuning methods like CoOp and VPT, we learn an incremental set of prompts, effectively mitigating the issue of incoming tasks overwriting previously learned parameters. Compared with prior CIL methods, it steers the learning of visual prompts using fixed textual prototypes, providing reliable anchor points in the embedding space. TPPT-VT can perform the best (and better than TPPT-V) consistently on the datasets with fine-grained classification characteristics, i.e., CUB, Aircraft, and Cars, which indicates that the textual prompt tuning helps to fill the gap on downstream tasks that have more specific characteristics. For the normal datasets, e.g. CIFAR100 and ImageNet-R, text prompt tuning may introduce training difficulties, influencing the performance of TPPT-VT.
Table 1: Average and final accuracy of different methods. All methods are trained with identical exemplar sizes, utilizing the CLIP model with a ViT-B/16 backbone using the same pre-trained weight from LAION-400M (Ilharco et al., 2021). $\bar{\mathcal{A}}$ denotes the average test accuracy over the incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and second best performances are indicated in red and blue, respectively.
@@ 4.2 Experimental Results
Comparisons with fine-tuning and adapter/LoRA-Based methods. In Table 2, we compare our method against fine-tuning-based approaches such as ZSCL (Zheng et al., 2023) and more recent adapter/LoRA (Hu et al., 2021)-based methods, including InfLoRA (Liang & Li, 2024) and MoE-Adapter (Yu et al., 2024). Despite utilizing additional reference data (Zheng et al., 2023), fine-tuning-based methods suffer from rapid forgetting due to full model updates and are computationally expensive. Adapter-based methods, on the other hand, struggle to maintain vision-text alignment across incremental learning stages, leading to suboptimal knowledge retention. In contrast, our approach leverages textual prototypes as stable anchors to guide the learning of incremental visual and multi-modal prompts. Our proposed method effectively mitigates representation drift, enhances cross-task consistency, and significantly improves continual learning performance while maintaining computational efficiency.
Table 2: Comparison with fine-tuning-based and adapter/LoRA-based methods on CIFAR-100 and Tiny-ImageNet using a 10-task split following the experiment settings of (Zheng et al., 2023; Yu et al., 2024). All methods utilize CLIP (ViT-B/16) with pre-trained weights from OpenAI (Radford et al., 2021).
@@ 4.3 Discussion
Learning with textual prototypes prevents representation drift. Catastrophic forgetting in continual learning is often attributed to representation drift (Gama et al., 2014; Lu et al., 2018), a phenomenon where the statistical properties of the target variable, which the model aims to predict, undergo changes during incremental learning. In Fig. LABEL:fig:drift, we measure representation drift during the incremental learning process by computing the divergence in the means of visual features for each class, relative to the class means from the previous incremental stage. We find that the phenomenon of representation drift (Gama et al., 2014; Lu et al., 2018) is severe, especially for the first few incremental stages. By enforcing visual features to be compact around their fixed textual prototypes through our proposed textual prototypical contrastive loss, we effectively mitigate catastrophic forgetting by significantly reducing representation drift.
Regularized textual prototypes encourage uniformity and prevent collapse. We quantify the pairwise distances between class means of visual features across different incremental stages in Fig. LABEL:fig:div. ZS-CLIP with pre-trained weight exhibits consistent uniformity (Wang & Isola, 2020; Cho et al., 2023) and is less prone to collapse (Jing et al., 2021; Bardes et al., 2021), due to the massive amount of data seen during pre-training. However, challenges arise when textual and visual prompts are learned jointly. In the initial incremental stages, the model tends to exhibit a tendency towards collapse, characterized by class means clustering too closely. The proposed diversity regularization effectively promotes uniformity of the model and prevents collapse.

### v3

@@ 3.2 Overview
We present the framework of our methods in Fig. 3. Previous CLIP-based CL methods (sprompts; attriclip; proof) usually treat the text encoder as a pre-trained classifier, overlooking the rich representation capability of the language modality. In Sec. 3.3 we introduce TPPT-V, which guides the learning of incremental visual prompts with static textual prototypes. Since learning with static textual prototypes may be limited by the quality of prototypes, in Sec. 3.4 we further propose TPPT-VT to learn textual prompts for higher quality and representative textual prototypes. Our approach encourages textual diversity to better regulate the learning process.
@@ 3.3 Incremental Visual Prompt Tuning with Textual Prototypes
Incremental visual prompt tuning. Prompt-based methods in CL have been proven to be highly effective in mitigating catastrophic forgetting, primarily by utilizing a fixed or incremental set of prompts (l2p; dualprompt; codaprompt). Building on this, we employ a straightforward method to prepend an incremental set of visual prompts on multiple Transformer layers (vpt; ivlp; promptsrc). These prompts are denoted as ${\mathbf{P}}_{v,l}=\{{\mathbf{P}}_{v,l}^{1},{\mathbf{P}}_{v,l}^{2},\cdots,{\mathbf{P}}_{v,l}^{M}\}$ for each Transformer layer $l$ ranging from the first layer up to the $d_{v}$-th layer of the vision encoder, where $M$ is the total number of prompts, and $d_{v}$ is a hyperparameter indicating the depth of Transformer layers at which prompts are learned.
@@
${\mathbf{P}}_{v,l}=\sum_{m=1}^{M}\alpha_{l}^{m}{\mathbf{P}}_{v,l}^{m},$ (2)
where the aggregation weights $\alpha_{l}=\text{LN}_{l}\big(q({\mathbf{x}})\big)\in\mathbb{R}^{M}$ are produced by a lightweight trainable linear layer applied at each layer with trainable prompts to obtain the affinities for the $M$ candidate prompts, in contrast to the complicated attended query-key matching with additional learnable attention vectors used in (codaprompt). Here, the query function $q(\cdot)=f_{\theta}(\cdot)\in\mathbb{R}^{D}$ refers to the pre-trained image encoder (l2p; dualprompt; codaprompt). Our concise formulation adopts the widely used query–key prompt selection/aggregation mechanism (l2p; dualprompt; codaprompt) and simplifies prior designs by omitting incremental learnable attention vectors and orthogonal regularization (codaprompt). The prompt tuning process is guided by consistent and stable knowledge derived from previously learned information through textual prototypes during incremental stages.
Anchoring the past with textual prototypes. Existing CLIP-based CIL methods (attriclip; continualclip) often utilize the text embeddings merely as a pre-trained classifier, thereby not fully exploiting the rich representational capacity of the language modality. To this end, we propose TPPT-V to guide the learning of incremental visual prompts through the use of static textual prototypes as anchor points in the embedding space. TPPT-V mitigates catastrophic forgetting by consistently ensuring that visual representations remain closely aligned and compact around corresponding static textual prototypes.
Specifically, our textual prototypes are derived from text embeddings, obtained through hand-crafted prompt templates with class names from OpenAI [^1] https://github.com/openai/CLIP whenever available. For datasets where prompt templates are not provided, we apply the basic template of “a photo of a [cls]”. To enhance the alignment (wang2020understanding) between visual features and their corresponding textual prototypes, we propose the Textual Prototypical Contrastive Loss (pcl; cmoa; li2023steering). This approach leverages the robust pre-trained representational capacity of these textual prototypes while preventing visual features from being assigned to the classes previously learned. It also effectively prevents the drift of previously learned semantics throughout the incremental learning phase (Fig. 4.2). We denote the Textual Prototypical Contrastive Loss as:
@@
where $\mathbbm{1}(\cdot)$ is the indicator function that selects samples corresponding to the given textual prototype. $\tilde{{\mathbf{z}}}=\tilde{f_{\theta}}({\mathbf{x}})$ is the visual embedding obtained through the prompted image encoder, and ${\mathbf{w}}_{c}$ is the text embedding for class $c$, acting as a static textual prototype. $N$ denotes the number of training examples in the current mini-batch, and $C$ is the number of classes already seen. At incremental step $t$, we construct the textual prototype set using only the class labels that have been introduced up to the current task. Text embeddings for future classes are not instantiated or used in any component.
@@
where $\mathcal{L}_{\text{CE}}$ represents the cross-entropy loss between predicted probabilities and ground truth. All loss terms are computed with the prompted image encoder $\tilde{f_{\theta}}$ and the frozen pre-trained text encoder $g_{\psi}$. This loss shares a similar form with the pre-training loss of the CLIP model (clip), but differs in purpose. While CLIP focuses on the one-to-one mapping between an image and its own caption to form a broad representation space, our method aims to cluster image embeddings to be compact around their corresponding class text embeddings, from the perspective of CL to mitigate forgetting.
@@ 3.4 Incremental multi-modal prompt tuning with regularized textual prototypes
Incremental textual prompt tuning. Although static textural prototypes from hand-crafted text prompts of classes effectively serve as stable anchors in TPPT-V, the effectiveness can be limited by the gap between pre-trained CLIP and specific downstream tasks (e.g., fine-grained classification), as well as the inherent discrepancy between the textual and visual embeddings in CLIP (liang2022mind). To address these limitations, we introduce TPPT-VT to further incorporate multi-model prompt tuning using learnable soft text prompts. Similar to the design of visual prompts in TPPT-V, we use an incremental set of textual prompts for each Transformer layer $l$, ranging from the first layer up to the $d_{t}$-th layer of the text encoder, where $d_{t}$ is set as a hyperparameter indicating the depth of Transformer layers at which prompts are learned. We denote these prompts as ${\mathbf{P}}_{t,l}=\{{\mathbf{P}}_{t,l}^{1},{\mathbf{P}}_{t,l}^{2},\cdots,{\mathbf{P}}_{t,l}^{C}\}$, where each class is assigned to a unique textual prompt, and $C$ represents the total number of classes encountered thus far. By using unique prompts for each class, TPPT-VT further enhances the fixed textual prototypes in TPPT-V, making them more representative.
Regularizing textual prompt tuning. While guiding the model with textual prototypes derived from hand-crafted prompts has demonstrated superior performance, the joint training of textual and visual prompts introduces new challenges. One critical issue is the model collapse, i.e, both branches of the joint-embedding model produce nearly identical and constant output vectors (jing2021understanding; vicreg) (as illustrated in Fig. 12(a)). This collapse compromises the model’s ability to distinguish between diverse inputs and effectively learn meaningful representations. In addition, naïve joint training can adversely affect the uniformity of the pre-trained model (wang2020understanding). Uniformity is a desirable property for joint-embedding models, ensuring that feature vectors are uniformly distributed roughly on the unit hypersphere. This uniform distribution is essential for preserving the maximum amount of information from the data.
To address these issues, we introduce the textual prototype diversity loss, which regularizes the model by explicitly enforcing textual prototypes to be uniformly distributed on the unit hypersphere. This is achieved by promoting high pairwise distances among textual prototypes (wang2020understanding; cho2023distribution), formally defined as follows:
@@ 4.1 Experiment Settings
Datasets. In line with common CIL benchmarks (icarl; l2p; dualprompt; codaprompt), we evaluate the performance using widely used datasets CIFAR100 (cifar), ImageNet-R (deng2009imagenet), TinyImageNet (le2015tiny), and CUB200 (WahCUB2002011). To further validate the efficacy of our approach, we extend our evaluation to include two additional datasets, FGVCAircraft (aircraft) and Stanford Cars (krause20133d), following the popular VLM methods (coop; kgcoop; maple; promptsrc). Specifically, we split the training datasets into 10 tasks, resulting in 10 classes per task for CIFAR100, Aircraft, and Stanford Cars, 20 classes per task for ImageNet-R, TinyImageNet, and CUB200. We sample 20 exemplars per class as our replay buffer as in (icarl).
Training details. For a fair comparison, we use the same backbone model CLIP with ViT-B/16, using pre-trained weights from (openclip; clip). We train all methods with a batch size of 64 for 5 epochs, using SGD with momentum for optimization. The learning rate starts from 0.1 and decays with cosine annealing. The visual prompt length $L_{v}$ and the textual prompt length $L_{t}$ are set to 4, and the depth for both visual prompts $d_{v}$ and textual prompts $d_{t}$ are set to 12. The number of prompts per task is set to 10. The loss weight $\alpha$ is set to 1 by default.
Evaluation metrics. To demonstrate the effectiveness of our approach, we report two commonly used metrics (icarl): (1) the final accuracy ${\mathcal{A}}_{T}$ after learning all the $T$ incremental stages, (2) the average accuracy $\bar{{\mathcal{A}}}=\frac{1}{T}\sum_{t=1}^{T}{\mathcal{A}}_{t}$ of the individual accuracies across the incremental stages.
Table 1: Average and final accuracy of different methods. All methods are trained with identical exemplar sizes, utilizing the CLIP model with a ViT-B/16 backbone using the same pre-trained weight from LAION-400M (openclip). $\bar{\mathcal{A}}$ denotes the average test accuracy over the incremental stages, $\mathcal{A}_{T}$ denotes the test accuracy at the last incremental stage. The best and second best performances are indicated in red and blue, respectively.
@@
(a) Experiment results across incremental stages. All methods are trained with the same exemplar size using the same pre-trained weights and backbone model of CLIP with ViT-B/16.
4.2 Experimental Results
Comparisons with prompt-based methods. In Table 1, we evaluate the final and average accuracies across various datasets, comparing current state-of-the-art methods. Additionally, we illustrate the performance across incremental stages in Fig. 8(a), and provide evaluations of multiple independent runs with varied orders of the incoming classes in Appendix C. Our proposed methods, TPPT-V and TPPT-VT, exhibit a remarkable improvement over previous approaches. In contrast to previous prompt tuning methods like CoOp and VPT, we learn an incremental set of prompts, effectively mitigating the issue of incoming tasks overwriting previously learned parameters. Compared with prior CIL methods, it steers the learning of visual prompts using fixed textual prototypes, providing reliable anchor points in the embedding space. TPPT-VT can perform the best (and better than TPPT-V) consistently on the datasets with fine-grained classification characteristics, i.e., CUB, Aircraft, and Cars, which indicates that the textual prompt tuning helps to fill the gap on downstream tasks that have more specific characteristics. For the normal datasets, e.g. CIFAR100 and ImageNet-R, text prompt tuning may introduce training difficulties, influencing the performance of TPPT-VT. In Table 2, we extend the comparisons to include two related methods DIKI and CPP, with further details in Appendix A.2.
Comparisons with fine-tuning and adapter/LoRA-Based methods. In Table 2, we compare our method against fine-tuning-based approaches such as ZSCL (zheng2023preventing) and more recent adapter/LoRA (lora)-based methods, including InfLoRA (inflora), MoE-Adapter (boosting), and CLAP4CLIP (jha2024clap4clip). Despite utilizing additional reference data (zheng2023preventing), fine-tuning-based methods suffer from rapid forgetting due to full model updates and are computationally expensive. Adapter-based methods, on the other hand, struggle to maintain vision-text alignment across incremental learning stages, leading to suboptimal knowledge retention. CLAP4CLIP adopts probabilistic fine-tuning with adapters and mixtures of task logits, which leaves the image–text embedding space under-constrained. In contrast, our approach leverages textual prototypes as stable anchors to guide the learning of incremental visual and multi-modal prompts. Our proposed method effectively mitigates representation drift, enhances cross-task consistency, and significantly improves continual learning performance while maintaining computational efficiency. Additional comparisons with the fine-tuning–based SPU method (zhang2024overcoming) are provided in Appendix Table 10.
Table 2: Comparison with prior methods on CIFAR-100 and Tiny-ImageNet using a 10-task split following the experiment settings of (zheng2023preventing; boosting). All methods utilize CLIP (ViT-B/16) with pre-trained weights from OpenAI (clip).
@@ 4.2 Experimental Results
(a) Representation drift $\downarrow$ (lower is better) across incremental stages for (a) CUB and (b) Aircraft. Representation drift measures the divergence in the means of visual features for each class compared to the last incremental stage.
(a) Pairwise distance $\uparrow$ (higher is better) between visual class means across incremental stages for (a) CUB and (b) Aircraft. A higher pairwise distance indicates the model has better learning capacity and is less prone to collapse.
@@ 4.3 Discussion
Learning with textual prototypes prevents representation drift. Catastrophic forgetting in continual learning is often attributed to representation drift (gama2014survey; lu2018learning), a phenomenon where the statistical properties of the target variable, which the model aims to predict, undergo changes during incremental learning. In Fig. 4.2, we measure representation drift during the incremental learning process by computing the divergence in the means of visual features for each class, relative to the class means from the previous incremental stage. We find that the phenomenon of representation drift (gama2014survey; lu2018learning) is severe, especially for the first few incremental stages. By enforcing visual features to be compact around their fixed textual prototypes through our proposed textual prototypical contrastive loss, we effectively mitigate catastrophic forgetting by significantly reducing representation drift.
Regularized textual prototypes encourage uniformity and prevent collapse. We quantify the pairwise distances between class means of visual features across different incremental stages in Fig. 12(a). ZS-CLIP with pre-trained weight exhibits consistent uniformity (wang2020understanding; cho2023distribution) and is less prone to collapse (jing2021understanding; vicreg), due to the massive amount of data seen during pre-training. However, challenges arise when textual and visual prompts are learned jointly. In the initial incremental stages, the model tends to exhibit a tendency towards collapse, characterized by class means clustering too closely. The proposed diversity regularization effectively promotes uniformity of the model and prevents collapse.